OpenStudy (anonymous):

dy/dt y=cos [sqrt(4t+11)]

OpenStudy (anonymous):

Chain rule.

OpenStudy (anonymous):

Let u = √4t+11) --> u' = ....

OpenStudy (anonymous):

I got the answer as \[-sinx \sqrt{4t+11} + \cos (2) (4t+11)^{-1/2}\]

OpenStudy (anonymous):

but the answer is actually with no (cos), what did I do wrong?

OpenStudy (anonymous):

Shouldn't be adding anything. Chain rule says to multiply all the derivatives together.

OpenStudy (anonymous):

( cosu)' = - u' sinu

OpenStudy (anonymous):

can you give me all steps so I can compare what I did wrong?

OpenStudy (anonymous):

There are three functions: cos [sqrt(4t+11)], sqrt(4t+11), and 4t+11. Find the derivatives of those three, then multiply them together.

OpenStudy (anonymous):

If it helps, think of it like this: y=cos W W=sqrt(U) U=4t+11

OpenStudy (anonymous):

it worked thanks

OpenStudy (anonymous):

so let's say if it's \[t ^{6}(t ^{5}+4)^{5}\] How would you divide out each individual like you just did

OpenStudy (anonymous):

Option one is to simplify by multiplying everything out then going term-by-term. Option two is to see that you have a product of U*V : U = t^6 V= W^5 W = t^5+4 You'll need a combination of product rule and chain rule here.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

how did they get the answer as \[t ^{5}(t ^{5}+4)^{4}(31^{5}+24)\]

OpenStudy (anonymous):

For \[\large \frac{d}{dx} t^6(t^5+4)^5 \space ?\]

OpenStudy (anonymous):

\[6t ^{5}(....)\]

OpenStudy (anonymous):

I don't understand this either, \[\large "t^5(t^5+4)^4(31^5+24)"\] What is that the answer to?

OpenStudy (anonymous):

to that same problem, but that is the answer in the solution manual

OpenStudy (anonymous):

Then I still don't get it.

OpenStudy (anonymous):

u = t^6 ---> u' = 6t^5 v = ( t^5 + 4) ^5 ---->v' = 5( t^5 + 4) ^4 * 5t^4 = 25 t^4 ( t^5 + 4) ^4

OpenStudy (anonymous):

,,,, = t^5 ( t^5 + 4) ^4 [ 6( t^5 + 4) + 25 t^5 ]

OpenStudy (anonymous):

= t^5 ( t^5 + 4) ^4 [ 31 t^5 + 24 ]

OpenStudy (anonymous):

So the textbook has typo at 31^5 !

OpenStudy (anonymous):

@thuyvy Questions?

OpenStudy (anonymous):

I got it, thanks :)

OpenStudy (anonymous):

you're the best chlorophyll, you saved me !! Thanks :)

OpenStudy (anonymous):

@Chlorophyll: can you help me on some other problems?

OpenStudy (anonymous):

Did you open new post, keep it simple one question per post will be less confused for the solver!